# FuseFormer: A Transformer for Visual and Thermal Image Fusion

### v1

@@ Abstract
Image fusion is the process of combining images from different sensors into a single image that incorporates all relevant information. The majority of state-of-the-art image fusion techniques use deep learning methods to extract meaningful features; however, they primarily integrate local features without considering the image’s broader context. To overcome this limitation, Transformer-based models have emerged as a promising solution, aiming to capture general context dependencies through attention mechanisms. Since there is no ground truth for image fusion, the loss functions are structured based on evaluation metrics, such as the structural similarity index measure (SSIM). By doing so, we create a bias towards the SSIM and, therefore, the input visual band image. The objective of this study is to propose a novel methodology for image fusion that mitigates the limitations associated with using evaluation metrics as loss functions. Our approach integrates a transformer-based multi-scale fusion strategy, which adeptly addresses both local and global context information. This integration not only refines the individual components of the image fusion process but also significantly enhances the overall efficacy of the method. Our proposed method follows a two-stage training approach, where an auto-encoder is initially trained to extract deep features at multiple scales at the first stage. For the second stage, we integrate our fusion block and change the loss function as mentioned. The multi-scale features are fused using a combination of Convolutional Neural Networks (CNNs) and Transformers. The CNNs are utilized to capture local features, while the Transformer handles the integration of general context features. Through extensive experiments on various benchmark datasets, our proposed method, along with the novel loss function definition, demonstrates superior performance compared to other competitive fusion algorithms. Furthermore, we show the effectiveness of the proposed fusion strategy with an ablation and performance analysis. The source code is also available. [^1] https://github.com/aytekXR/FuseFormer-Infrared-Fusion
@@
Image Fusion, Visual Infrared Image Fusion, Transformer Based Image Fusion, Structural Similarity Metric
@@ I Introduction
Image fusion is a powerful technique in the context of computer vision that involves combining information from multiple images taken at different wavelengths or bands to create a single, comprehensive representation. The primary objective of image fusion is to extract and integrate complementary details and features from each input image, resulting in a more informative and enhanced composite image. The fusion of images from different bands holds immense significance and due to fact that its widely used in night vision and thermal imaging, remote sensing and multiple spectre imaging, last but not least multi-modal medical imaging.
(a) Visual Band, Thermal Band Images
(b) Our Output Image, SwinFusion[1] Image
(c) M3FD[2] - RFN-Nest[3] Output Images
(d) IFT[4], DenseFuse[5] Output Images
Figure 1: Night Vision Enhancement Comparison with SoTA Method
Figure 2: Proposed Method’s Block Diagram
Research on image fusion dates back to 1989 with Toet et al. [6]. Traditional image fusion algorithms have been extensively investigated in the existing literature, but they are not without drawbacks. A significant concern with these methods lies in the incorporation of handcrafted steps, leading to suboptimal outcomes. With the advent of the deep learning era, attempts have been made to address this issue using Convolutional Neural Networks (CNNs) [7], Autoencoders [5, 8, 9, 10, 11, 12], and Generative Adversarial Networks (GANs) [13, 14, 15, 16, 17, 18, 19, 20]. However, these approaches exhibit limitations in effectively capturing long-range dependencies. To mitigate this challenge, Transformers [21, 22, 23, 4, 24, 1, 25, 26, 1, 27] have been employed to tackle image fusion problems. Despite this advancement, a primary drawback persists: the reliance on evaluation metrics such as Structural Similarity Index (SSIM) [28] as the main component of the loss function. This practice, while yielding satisfactory quantitative results, often leads to poor qualitative outcomes in the fused images. The main contributions of this work can be summarized as follows:
  • We propose a novel fusion method, called FuseFormer, that utilizes a Transformer-CNN fusion block with a unique loss function by taking both input images into account. By doing so, it mitigates the gap between quantitative and qualitative results.
  • The proposed method utilizes Transformers to capture global context and combines the results with local features from Convolutional Neural Networks (CNNs).
@@ II Related Work
The RGB and infrared image fusion domain has undergone extensive research, traversing from Traditional Fusion Algorithms to cutting-edge Transformer-based models [29, 30, 31, 32, 33]. In the early 1990s, methods like Sparse Representation and Multi-scale Transformation were explored, each with its inherent limitations. Traditional algorithms, relying on handcrafted steps, faced challenges in adaptability and time complexity [29, 30, 31, 32, 33]. The scarcity of labeled datasets for RGB-IR fusion prompted a shift towards unsupervised scenarios, guiding our exploration of Performance Evaluation metrics.
With the advent of deep learning, learning-based algorithms became predominant, categorized by learning methods, loss functions, and the use of labeled datasets [34, 5, 8, 9, 13]. CNN-based approaches, both supervised and unsupervised, exhibited success in feature extraction for image fusion, yet challenges persisted in scenarios with significant differences in factors like illumination or resolution [34]. Autoencoder-based algorithms, utilizing neural networks for dimensionality reduction, showcased advancements in works such as DenseFuse, Raza et al., and Fu et al. [5, 8, 9].
GAN-based methods, introduced by Goodfellow et al., focused on unsupervised fusion, integrating attention mechanisms and residual connections for improved performance [13, 18, 35]. While these approaches demonstrated promise, challenges persisted in effectively handling the inherent differences between fused and source images.
A transformative shift in RGB-IR image fusion occurred with the introduction of Transformer-based algorithms in 2021 [21, 22, 23]. These methodologies, driven by the self-attention mechanism, marked a paradigm shift by efficiently managing long-range dependencies in images. Innovative designs, such as multiscale fusion strategies and dual transformer approaches, were introduced, emphasizing the seamless integration of Transformers with traditional methods [4, 26, 24, 36]. Unsupervised Transformer-based techniques, reliant on loss functions, eliminated the need for labeled data but posed challenges in methodological evaluations [4, 26]. Ongoing research explores diverse Transformer integrations, Transformer-CNN combinations, and the utilization of auxiliary information to further enrich the fusion process [4, 26, 24, 36]. The integration of deep learning methods has not only enhanced feature extraction capabilities but has also paved the way for more adaptive and robust solutions in the challenging methodologies of image fusion. Due to the vastness of the subject, it is impractical to encompass all relevant studies here. For a comprehensive overview, readers are directed to Zhang et al. [37] which provides insights into the broader landscape of RGB-IR image fusion research, including advancements, methodologies, and challenges.
@@ III Methodology
Image fusion can be decomposed into three key components: the feature extractor, feature fuser, and image reconstructor. The feature extractor is responsible for extracting multilevel features from input images. Subsequently, the feature fuser merges these extracted features into unified feature maps for each level. These consolidated features play a crucial role in the final image reconstruction, orchestrated by the image reconstructor.
In terms of analogy, the feature extractor corresponds to the encoder in autoencoders, while the image reconstructor aligns with the decoder in autoencoders. The feature fuser leverages both Convolutional Neural Networks (CNNs) and transformers, with CNNs adeptly managing local features and transformers overseeing global contexts. This merge enhances the fusion process, ultimately aiming to improve accuracy.
Our approach involves a two-stage training process. In the initial stage, we train an autoencoder, originally derived from RFN-Nest [3] as illustrated in Figure 3. Subsequently, in the second stage, we train our fusion block, depicted in Figure 4, in conjunction with the previously trained autoencoder.
@@ III-A Autoencoder Training
(a) Block Diagram
(b) RFN-Nest [3] Encoder
(c) RFN-Nest [3] Decoder
Figure 3: Autoencoder Training
The initial phase of the training process involves instructing the encoder network to capture multi-scale deep features. Concurrently, the decoder network is also trained to reconstruct the input image, utilizing the aforementioned multi-scale deep features. The training framework of the auto-encoder network is depicted in Figure 3. These extracted multi-scale deep features are then fed into the decoder network for the purpose of reconstructing the input image. Leveraging short cross-layer connections ensures the comprehensive utilization of the multi-scale deep features in the image reconstruction process.
The loss function, denoted as $L_{ae}$, serves as the training criterion for the autoencoder network and is defined in the subsequent manner:
@@
$L_{\text{pixel}}=\left\lvert\left\lvert\text{image}_{\text{output}}-\text{image}_{\text{input}}\right\rvert\right\rvert_{F}^{2}$ (2)
@@
The second term in Eq 1 is the SSIM loss $L_{SSIM}$ is defined as in Eq 3:
$L_{SSIM}=1-SSIM(image_{output},image_{input})$ (3)
where $SSIM(.)$ is the structural similarity measure [28] which quantifies the structural similarity of the two images. The structural similarity between Input and Output is constrained by $L_{SSIM}$. The Structural Similarity Index (SSIM) is a widely used metric for evaluating the similarity between two images. It aims to capture not only the pixel-wise differences but also the structural information and perceptual quality of the images.
The Structural Similarity Index ($SSIM$) is a metric that quantifies the similarity between two images, yielding values within the range of -1 to 1. A $SSIM(\bigodot,\bigodot)$ value of 1 denotes perfect similarity, indicating that the images share same characteristics in terms of luminance, contrast, and structure. Conversely, a value close to -1 signifies a substantial dissimilarity between the images. Notably, the $SSIM(\bigodot,\bigodot)$ index demonstrates a strong correlation with human perception of image quality, making it widely employed in diverse image processing and computer vision applications [28].
The $SSIM$ constrains its output to the range of $[-1,1]$, which consequently bounds the $L_{SSIM}$ loss function (as defined in Eq. (3)) to the interval $[0,2]$. In this context, lower values of $L_{SSIM}$ indicate better performance with respect to $SSIM$. In contrast, the $L_{pixel}$ loss is unbounded. To balance the impact of both $L_{pixel}$ and $L_{SSIM}$ during training, the trade-off parameter $\alpha$ in Eq. (1) governs their relative magnitudes.
In short, the autoencoder depicted in Figure 3 is subjected to training using the MS-COCO dataset [38] and the RoadScene dataset [39]. The training process is guided by the loss function presented in Eq. (1).
@@ III-B Fusion Block Training
(a) Block Diagram
(b) Fusion Block
Figure 4: Fusion Block Training
After successfully extracting multi-scale multi-level features in first stage, see Figure 3, the aim in the second stage training is to merge the features from two separate images: visible band and infrared band images. In this stage, the outputs from separate encoders are then merged into a single multi-scale feature map with fusion block depicted in Figure 4 per scale. This fusion process, which lays on the hearth of this research, is a crucial step as it combines diverse features from different bands, enhancing the feature representation. It allows the model to leverage the strengths of each band, thereby improving the overall effectiveness of the feature extraction process.
Following the fusion, the resulting multi-scale feature maps are subsequently decoded, leading to the reconstruction of the original image, as depicted in Figure 4. Decoding phase reconstructs an image that encapsulates the combined information from both bands. In essence, the reconstructed image, though visually similar to the input images, carries a much richer set of features, potentially paving the way for more accurate subsequent analyses or processes.
In the fusion block, Conventional Convolutional Neural Network (CNN) based techniques facilitate image fusion through the merge of local features. However, a significant limitation inherent to these methods is their lack of consideration for the global context and long range dependencies. In an attempt to overcome this limitation, transformer-based models have been introduced, which leverage the attention mechanism to effectively model the global context.
The first stage of this training protocol necessitates the use of an auto-encoder to extract deep, multi-scale features. In the subsequent stage, these multi-scale features are blended via a fusion strategy that innovatively combines CNNs with Transformers. Comprising a CNN and a transformer branch, the combined fusion blocks capably capture both local and global context features. This method’s success lies in its ability to leverage the strengths of both CNNs and Transformer models, providing a comprehensive view of an image by capturing both local and global contexts. The strategy’s robustness, combined with its efficiency, heralds a new direction for further developments in the field of image fusion. Used combined fusion block details can be seen at Figure 3(b).
The fusion block is characterised by its dual-branch design, which consists of a spatial branch and a transformer branch. The spatial branch integrates convolution layers and a bottleneck layer, specifically tailored to distil local feature representations. Conversely, the transformer branch employs an axial attention-based transformer block to capture the global context embedded within the input data.
While designing the loss function, current researches continue with the same loss defined as in Eq 1. Only single input image, visible band image in general, and the output fused image is used in this loss function. Skipping the unused input image, infrared band image in general, yields a overfitting towards to the used input image. It can be explained as:
@@
  • $L_{ssim}=0$ if and only if the input and output image is identical by definition of $SSIM(\bigodot,\bigodot)$.
Therefore, a unique loss function is deployed that leverages both input images. The fusion loss function $L_{fuse}$, as in $L_{ae}$, can be formulated as in Eq 4:
$L_{fuse}=L_{pixel}+\alpha L_{ssim}$ (4)
This fusion loss function aims to balance the contribution from pixel-level losses, denoted as $L_{feature}$, and structural similarity losses, $L_{ssim}$, modulated by a trade-off factor, $\alpha$ which handles the difference of order of magnitude for losses. To ensure $L_{fuse}=0$ in Eq 4 corresponds to an optimal fusion scenario, the definitions of $L_{ssim}$ and $L_{feature}$ must be updated as shown in Eq 6 and Eq 5 respectively. While re-defining the losses, the following constraints need to be considered:
@@
    The fused image should have a higher resemblance to the visual band image while maintaining the global context from the infrared band image almost identical to the visual band image. As a result, $L_{ssim}$ must be computed for both input visual and infrared band images, ideally but not necessarily favoring the visual band image.
@@
The SSIM loss $L_{ssim}$ is then redefined as:
$L_{ssim}=\left[1-SSIM(I_{f},I_{v})\right]^{2}+\left[1-SSIM(I_{f},I_{i})\right]^{2}$ (5)
The redefined $L_{ssim}$ is capable of measuring the similarity of the fused image to both visual and infrared images, and is limited to the interval $(0,8]$.
The pixel-wise loss $L_{feature}$ can be formulated as:
@@
Here, $M$ refers to the number of scales for deep feature extraction, while $f$, $vi$, and $ir$ denote the fused image, the input visual band image, and the input infrared band image respectively. $\omega^{m}$, $\omega_{vi}$, and $\omega_{ir}$ represent trade-off parameters employed to harmonize the magnitudes of the losses. $\phi_{x}^{m}$ corresponds to the feature maps of $imagex$, which could be either the input or output feature maps of the fusion block, as depicted in Figure 4.
@@ IV Results
For the hardware components, an NVIDIA RTX 3060Ti with 16GB of memory, paired with an Intel i9 10th generation CPU, was utilized. During the comparison of models, the following metrics were widely used for the evaluation: Entropy (En)[42],Sum of the Correlations of Differences (SCD)[43], Mutual Information (MI)[44],Structural Similarity Index Metric (SSIM)[28].
TABLE I: Loss Functions Comparison
Method Entropy[42]normal-$\uparrow$ SCD[43]normal-$\downarrow$ MI[44]normal-$\uparrow$ SSIM[28]normal-$\uparrow$
Proposed $L_{fuse}$ in Eq 4 4.536 5.433 1.591 0.884
$L_{ae}$ as $L_{fuse}$ Exp 4.559 6.466 0.552 0.879
RFN-Nest[3] 4.729 7.062 0.602 0.541
(a) Visual Band Images
(b) Infrared Band Images
(c) Model With New $L_{fuse}$ Loss Output Images
(d) Model With $L_{fuse}=L_{ae}$ Loss Output Images
Figure 5: Loss Functions Comparison
In Table I, the impact of redefining the loss function is systematically examined. Reference results from RFN-Nest [3] are provided since the autoencoder is inherited from RFN-Nest. Despite considering not only the visible band image in the redefined $L_{ssim}$, as depicted in Eq 5, the resulting structural similarity index ($SSIM$) of the visible band image and the fused image surpasses that of the original loss function component $L_{ssim}$, detailed in Eq 3. The redefinition of $L_{ssim}$ not only enhances quantitative results but also improves qualitative outcomes, as illustrated in Figure 5.
(a) Visual Band Images
(b) Infrared Band Images
(c) Our Output Images
(d) SwinFusion[1] Output Images
Figure 6: Comparison with SoTA
(a) M3FD[2] Output Images
(b) IFT[4] Output Images
(c) RFN-Nest[3] Output Images
(d) DenseFuse[5] Output Images
Figure 7: Comparison with SoTA cont’d
TABLE II: Comparison with SoTA
Method Entropy[42]normal-$\uparrow$ SCD[43]normal-$\downarrow$ MI[44]normal-$\uparrow$ SSIM[28]normal-$\uparrow$
Ours 4.536 5.433 1.591 0.884
SwinFusion[1] 4.605 6.760 0.804 0.690
M3FD[2] 4.625 6.858 0.742 0.659
IFT[4] 4.644 6.864 0.684 0.630
DenseFuse[5] 4.724 6.455 0.853 0.588
RFN-Nest[3] 4.729 7.062 0.602 0.541
In Table II, FuseFormer is systematically compared with state-of-the-art (SoTA) methods in a quantitative manner. Simultaneously, a qualitative comparison is conducted through the examination of Figures 6 and 7. The ordering of methods in Table 5 is based on the structural similarity index (SSIM) [28]. Despite variations in the claims made by the authors of the referenced papers, the utilization of the TNO inference approach resulted in the compilation presented in Table II.
The comprehensive evaluation, encompassing both Table II and Figures 6 and 7, establishes FuseFormer as superior both qualitatively and quantitatively across nearly all metrics.
(a) Visual Band Images
(b) Infrared Band Images
(c) Our Output Images
(d) SwinFusion[1]Output Images
Figure 8: Night Vision Enhancement
(a) M3FD[2] Output Images
(b) IFT[4] Output Images
(c) RFN-Nest[3] Output Images
(d) DenseFuse[5] Output Images
Figure 9: Night Vision Enhancement cont’d
As a practical example of using infrared-visible band images, we selected suitable images from the TNO dataset for night vision applications. In Figures 8 and 8, the input visible and infrared images represent instances of night vision scenarios. This specific example is chosen to effectively demonstrate the long-range dependencies and global context present in such images.The images in the second and third columns predominantly depict scenes captured under low-light conditions. Although, at a quick qualitative glance, the differences between the images may seem subtle, a closer examination reveals nuanced distinctions. Turning our attention to Table I, a thorough evaluation indicates that our proposed method consistently produces superior results. In comparison to state-of-the-art (SoTA) techniques, our method shows significant improvements across nearly every evaluated performance metric.
TABLE III: Model Tuning Results: Improvement Rate

### v2

@@ Abstract
Due to the lack of a definitive ground truth for the image fusion problem, the loss functions are structured based on evaluation metrics, such as the structural similarity index measure (SSIM). However, in doing so, a bias is introduced toward the SSIM and, consequently, the input visual band image. The objective of this study is to propose a novel methodology for the image fusion problem that mitigates the limitations associated with using classical evaluation metrics as loss functions. Our approach integrates a transformer-based multi-scale fusion strategy that adeptly addresses local and global context information. This integration not only refines the individual components of the image fusion process but also significantly enhances the overall efficacy of the method. Our proposed method follows a two-stage training approach, where an auto-encoder is initially trained to extract deep features at multiple scales in the first stage. For the second stage, we integrate our fusion block and change the loss function as mentioned. The multi-scale features are fused using a combination of Convolutional Neural Networks (CNNs) and Transformers. The CNNs are utilized to capture local features, while the Transformer handles the integration of general context features. Through extensive experiments on various benchmark datasets, our proposed method, along with the novel loss function definition, demonstrates superior performance compared to other competitive fusion algorithms.
@@
RGB-T Image Fusion, Vision Transformers, Structural Similarity Metric
@@ I Introduction
Image fusion is a powerful technique in the context of computer vision that involves combining information from multiple images taken at different wavelengths or bands to create a single, unified representation. The primary objective of image fusion is to extract and integrate complementary details and features from each input image, resulting in a more informative and enhanced composite image. The fusion of images from different bands is widely used in various applications, such as night vision [1] and thermal imaging [2], remote sensing [3] and multi-modal medical imaging [4], just to name a few.
(a) Visual band
(b) Thermal band
(c) FuseFormer
Figure 1: Example output from the proposed FuseFormer model: Fusion of visual and thermal images.”
Research on image fusion dates back decades, with [5] being one of the pioneering studies. Traditional image fusion algorithms—that is, those prior to deep learning—have been thoroughly studied in the literature [6]. A significant concern with these methods lies in the incorporation of handcrafted steps, leading to suboptimal outcomes for changing conditions. With the deep learning (DL) era, attempts have been made to address this issue using Convolutional Neural Networks (CNNs) [7, 8, 9, 10], autoencoders [11, 12], attention models [13], and Generative Adversarial Networks (GANs) [14, 15, 16, 17, 18, 19, 20, 21]. These first-generation of DL approaches are significantly better than the pre-DL era methods, but still exhibit limitations in effectively capturing long-range dependencies (i.e. global context) within the image. To address this challenge, Vision Transformers [22, 23] have been utilized for image fusion tasks [24, 25, 26, 27, 28, 29, 30].
Despite these significant advancements, a notable gap is apparent in the literature: the reliance on evaluation metrics such as the Structural Similarity Index (SSIM) [31] as the primary component of the loss function. While this approach may produce satisfactory quantitative results, it may provide nonoptimal results [32]. Our contributions to this issue, as well as other aspects of our work, can be summarized as follows:
  • We propose a novel fusion framework, called the FuseFormer, that utilizes a Transformer+CNN fusion block with a unique loss function that takes both modalities into account. By doing so, it mitigates the gap between quantitative and qualitative results.
  • The proposed method utilizes Transformers to capture global context and combines the results with local features obtained from Convolutional Neural Networks (CNNs).
@@ II Related Work
As briefly introduced in the previous section, in the last four decades of research, the RGB and infrared image fusion domain has been the subject of extensive research, spanning from traditional fusion algorithms [6] to state-of-the-art Transformer-based models [33, 34, 35, 36, 37]. Traditional algorithms, relying on handcrafted steps, faced challenges in adaptability and time complexity [33, 34, 35, 36, 37]. The scarcity of labeled datasets for RGB-IR fusion prompted a shift towards unsupervised scenarios [38], which also guided our investigation towards enhancing evaluation metrics in this paper.
With the advent of deep learning, learning-based algorithms became predominant, categorized by learning methods, loss functions, and the use of labeled datasets [39, 8, 9, 14]. CNN-based approaches, both supervised and unsupervised, exhibited success in feature extraction for image fusion, yet challenges persisted in scenarios with significant changes in factors like illumination or resolution [39]. Autoencoder-based algorithms, utilizing neural networks for dimensionality reduction, showcased advancements in works such as [12, 11].
GAN-based methods focused on unsupervised fusion, integrating attention mechanisms and residual connections for improved performance [14, 19, 40]. While these approaches demonstrated promise, challenges persisted in effectively handling the inherent differences between fused and source images.
A literally “transformative” shift in RGB-IR image fusion occurred with the introduction of Vision Transformer-based algorithms in 2021 [22, 23, 24]. These methodologies, driven by the self-attention mechanism, marked a paradigm shift by efficiently managing long-range dependencies in images. Innovative designs, such as multiscale fusion strategies and dual transformer approaches, were introduced, emphasizing the seamless integration of Transformers with traditional methods [25, 29, 26, 41]. Unsupervised Transformer-based techniques, reliant on loss functions, eliminated the need for labeled data but posed challenges in methodological evaluations [25, 29]. Ongoing research explores diverse Transformer integrations, Transformer-CNN combinations, and the utilization of auxiliary information to further enrich the fusion process [25, 29, 26, 41]. The integration of deep learning methods has not only enhanced feature extraction capabilities but has also paved the way for more adaptive and robust solutions in the challenging methodologies of image fusion. For a comprehensive overview, readers are directed to Zhang et al. [42] which provides insights into the broader landscape of RGB-IR image fusion research, including advancements, methodologies, and challenges. For a review of RGB-IR image and videos set,which serves as the application data for fusion algorithms, readers may refer to [43].
@@ III Methodology
In general terms, the image fusion operation can be decomposed into three key components: the feature extractor, feature fuser, and the image reconstructor. Although these components may differ from architecture to architecture, it is possible to define their purposes such that the feature extractor is responsible for extracting multilevel features from input images. Subsequently, the feature fuser merges these extracted features into unified feature maps for each level. These consolidated features play a crucial role in the final image reconstruction, orchestrated by the image reconstructor.
Figure 2: FuseFormer Block Diagram
The same idea may be used to theoretically explain the model architecture that we present in this article as seen in Figure 2. In terms of analogy, the feature extractor corresponds to an encoder, while the image reconstructor aligns with a decoder for modern DL architectures. Both the visible and IR-band images are fed into this encoder first. The encoder outputs multiscale features, denoted as $\phi_{Vi}$ and $\phi_{IRi}$. These features are then fed into the proposed a feature fuser (the middle component in Figure 2), which leverages both Convolutional Neural Networks (CNNs) and transformers, to manage both local features and global contexts. In the following, we provide the architectural details and the strategies we use to separately train each of these components.
@@ III-A Autoencoder Training
Our approach involves a two-stage training process. In the initial stage, we concurrently train the encoder and the decoder in Figure 2, by training an autoencoder, originally derived from RFN-Nest [44] as illustrated in Figure 3. Subsequently, in the second stage, we train our fusion block, depicted in Figure 4, in conjunction with the previously trained encoder and the decoder.
Figure 3: The autoencoder architecture, derived from RFN-Nest [44]
The initial phase of the training process involves instructing the encoder network to capture multi-scale deep features. Concurrently, the decoder network is also trained to reconstruct the encoder output. The extracted multi-scale deep features ($\phi_{IRi}$ and $\phi_{Vi}$) are fed into the decoder network to reconstruct the input image. As explained in detail in [44], leveraging short cross-layer connections ensures the utilization of the multi-scale deep features in the image reconstruction process.
While training the autoencoder, the loss function, denoted as $L_{ae}$, serves as the main criterion and is defined in [44] in the subsequent manner:
@@
$L_{\text{pixel}}=\left\lvert\left\lvert\text{I}_{\text{output}}-\text{I}_{\text{input}}\right\rvert\right\rvert_{F}^{2}$ (2)
@@
The second term in Eq 1 is the SSIM loss $L_{SSIM}$ and is defined as:
$L_{SSIM}=1-SSIM(I_{output},I_{input})$ (3)
where $SSIM(.)$ is the structural similarity measure [31] which quantifies a measure of similarity between the two images. $L_{SSIM}$. SSIM is a widely used metric for evaluating the similarity between two images. It aims to capture not only the pixel-wise differences but also the structural information and perceptual quality of the images. ($SSIM$) metric outputs values within the range of -1 to 1. The output value 1 for the $SSIM(\cdot,\cdot)$ function denotes perfect similarity, indicating that the images share same characteristics in terms of luminance, contrast, and structure. Conversely, a value close to -1 signifies a substantial dissimilarity between the images. Notably, the $SSIM(\cdot,\cdot)$ index demonstrates a strong correlation with human perception of image quality, making it widely employed in diverse image processing and computer vision applications [31].
The $SSIM$ constrains its output to the range of $[-1,1]$, which consequently bounds the $L_{SSIM}$ loss function (as defined in Eq. 3) to the interval $[0,2]$. In this context, lower values of $L_{SSIM}$ indicate better performance with respect to $SSIM$. In contrast, the $L_{pixel}$ loss is unbounded. To balance the impact of both $L_{pixel}$ and $L_{SSIM}$ during training, the trade-off parameter $\alpha$ in Eq. 1 governs their relative magnitudes.
@@ III-B Fusion Block Training
Figure 4: The Fusion Block
After successfully extracting multi-scale multi-level features in first stage, the aim in the second stage training is to merge the features from the two input images. In this stage, the RGB and IR features extracted from the encoder are merged into a single multi-scale feature map with fusion block as depicted in Figure 4. This fusion process aims at combining diverse features from different bands. The fusion block is characterised by its dual-branch design, which consists of a spatial branch and a transformer branch. The spatial branch consisting of CNN blocks facilitates fusion through the merge of local features. CNNs provide regional spatial coding but they lack the ability to model the global context and long spatial dependencies. To overcome this limitation, transformer-based models are included in the Fusion Block, which utilize axial attention mechanism to effectively model the global context. As can be seen in Figure 4, for each visible and infrared band image, two multi-head axial attention blocks are utilized, separately for each image dimension. The feature output of the attention blocks is concatenated with the original encoder features (i.e., $\phi_{IRi}$ and $\phi_{Vi}$) before being convolved for the final $\phi_{Fi}$ fused image construction.
Following the fusion, the resulting multi-scale feature maps, $\phi_{Fi}$, are subsequently decoded, leading to the reconstruction of the original image. Decoding phase reconstructs an image that encapsulates the combined information from both bands. Hence, the reconstructed image, though visually similar to the input images, carries a richer set of features.
III-C Fusion Loss
In the process of designing the loss function, contemporary studies persist with the identical loss formulation as presented in Eq. 1. This loss function incorporates a single input image, typically the visible band image, along with the output fused image. By neglecting the unused input image, typically the infrared band image, there arises a risk of overfitting towards the utilized input image. We can examine the edge loss circumstances in order to better visualize this phenomenon:
@@ III-C Fusion Loss
  • $L_{ssim}=0$ if and only if the input and output image is identical by definition of $SSIM(\cdot,\cdot)$.
A bias toward the SSIM and consequently the visible band picture is unavoidably created when the loss function is built as a combination of the two above. To overcome this, we propose a unique loss function. Namely, the fusion loss function $L_{fuse}$, similarly to $L_{ae}$, can be formulated as in Eq 4:
$L_{fuse}=L_{feature}+\alpha L_{\overline{ssim}}$ (4)
This fusion loss function aims to balance the contribution from pixel-level losses, denoted as $L_{feature}$, and structural similarity losses, $L_{\overline{ssim}}$, modulated by a trade-off factor, $\alpha$ which handles the difference of order of magnitude for losses. While defining these two loss components, the following constraints need to be considered:
@@
    The fused image should have a higher resemblance to the visual band image while maintaining the global context from the infrared band image almost identical to the visual band image. As a result, $L_{\overline{ssim}}$ must be computed for both input visual and infrared band images, ideally but not necessarily favoring the visual band image.
@@
Then, The updated SSIM loss, $L_{\overline{ssim}}$, is defined as:
$L_{\overline{ssim}}=\left[1-SSIM(I_{f},I_{v})\right]^{2}+\left[1-SSIM(I_{f},I_{i})\right]^{2}$ (5)
The updated $L_{\overline{ssim}}$ is capable of measuring the similarity of the fused image to both visual and infrared images, and is limited to the interval $(0,8]$.
The updated pixel-wise loss $L_{feature}$ can be formulated as:
@@
Here, $M$ refers to the number of scales for deep feature extraction, while $f$, $vi$, and $ir$ denote the fused image, the input visual band image, and the input infrared band image respectively. $\omega^{m}$, $\omega_{vi}$, and $\omega_{ir}$ represent trade-off parameters employed to harmonize the magnitudes of the losses. $\phi_{f}^{m}$ corresponds to the feature maps of an image scale $m$, as depicted in Figure 2.
@@
The hardware setup comprised an NVIDIA RTX 3060Ti featuring 16GB of memory, combined with an Intel i9 10th generation CPU. Throughout the model comparison, the assessment primarily relied on the following metrics: Entropy (En)[48], Sum of the Correlations of Differences (SCD)[49], Mutual Information (MI)[50], and Structural Similarity Index Metric (SSIM)[31].
IV Experiments and Results
TABLE I: Quantitative comparison of the loss Functions
Method Entropy[48$\uparrow$ SCD[49$\downarrow$ MI[50$\uparrow$ SSIM [31$\uparrow$
$L_{fuse}$ 4.536 5.433 1.591 0.884
$L_{ae}$ 4.559 6.466 0.552 0.879
RFN-Nest 4.729 7.062 0.602 0.541
Our initial set of experiments focuses on analyzing the impact of the updated loss function. We primarily train the proposed FuseFormer architecture separately using the loss functions $L_{fuse}$ and $L_{ae}$. In addition, the results of the autoencoder architecture (without the fusion block) and using the conventional loss function $L_{ae}$, are reported from the RFN-Nest [44] original paper. In Table I, comparative results with different evaluation metrics, including Entropy [48], SCD [49], MI [50], and SSIM [31] are provided. The proposed FuseFormer architecture, trained using the $L_{fuse}$ loss function performes the best in 3 of the 4 evaluation metrics, including the original SSIM metric, although not only the visible band image is utilized for the updated SSIM loss metric $L_{\overline{SSIM}}$. When we observe samples as seen in Figure 5, we can see that redefinition of $L_{\overline{SSIM}}$ not only enhances quantitative results but also improves qualitative outcomes. For instance, in the leftmost image, the person laying on the ground shows better visibility in Figure 5(c), when compared to Figure 5(d).
(a) Visual-band Images
(b) Infrared Images
(c) FuseFormer trained with $L_{fuse}$ Loss
(d) FuseFormer trained with $L_{ae}$ Loss
Figure 5: Visual comparison of the loss functions
TABLE II: Quantitative comparison against the SoTa methods.
Method Entr.[48$\uparrow$ SCD[49$\downarrow$ MI[50$\uparrow$ SSIM [31$\uparrow$
FuseFormer 4.536 5.433 1.591 0.884
SwinFusion 4.605 6.760 0.804 0.690
M3FD 4.625 6.858 0.742 0.659
DenseFuse 4.724 6.455 0.853 0.588
RFN-Nest 4.729 7.062 0.602 0.541
In our second set of experiments, we benchmark the FuseFormer architecture trained with the proposed $L_{fuse}$ loss function, against the state-of-the-art (SoTa) fusion methods from the literature. For this comparison, we utilize SwinFusion[27], M3FD[51], RFN-Nest[44] and DenseFuse[8] models. In Table II, quantitative results are provided. In parallel with the result of Table I, FuseFormer performs the best for the same three evaluation metrics, including the original SSIM. In Figure 6, a qualitative comparison is also depicted.
For the purpose of doing a visual analysis for night vision settings, we carefully select a group of images from the TNO dataset that primarily show scenes taken in low light. These images are the two right-most columns in Figure 6. In this figüre, the original visible-band and infrared images, our results, and the result of the selected SoTa methods are provided. Lighting at night creates a global effect in these images, hence, they effectively demonstrate the long-range dependencies (i.e. global context) present in these scenarios. We observe that FuseFormer outperforms or is comparable with the SoTa approaches, in all scenarios.
TABLE III: Model tuning Experiments: Transformer Layers
Exp. Entropy[48$\uparrow$ SCD[49$\downarrow$ MI[50$\uparrow$ SSIM [31$\uparrow$
$T_{12}$ 4.533 6.152 1.346 0.846
$T_{10}$ 4.535 6.335 1.232 0.824
$T_{8}$ 4.587 6.657 0.874 0.719
TABLE IV: Model tuning Experiments: Batch Size
Exp. Entropy[48$\uparrow$ SCD[49$\downarrow$ MI[50$\uparrow$ SSIM [31$\uparrow$
$B_{4}$ 4.542 6.478 1.130 0.746
$B_{2}$ 4.587 6.657 0.874 0.719
TABLE V: Model tuning Experiments: Learning Rate
Exp. Entropy[48$\uparrow$ SCD[49$\downarrow$ MI[50$\uparrow$ SSIM [31$\uparrow$
$LR_{1e-6}$ 4.553 6.551 1.032 0.774
$LR_{1e-5}$ 4.570 6.603 0.948 0.747
$LR_{1e-4}$ 4.587 6.657 0.874 0.719
TABLE VI: Model tuning Experiments: Improvement Rate